# Towards Better Benchmark Datasets for Inductive Knowledge Graph Completion

### v1

@@ 4.2 Analysis of New Datasets
In this section we analyze the new inductive datasets created following the partition-based procedure outlined in Section 4.1. We create (E) inductive datasets, from WN18RR [26], CoDEx-M [31], and HetioNet [32]. For the (E, R) datasets, we use CoDEx-M [31] and FB15k-237 [28]. Each dataset has two separate inference graph except for CoDEx-M on the (E) task which has one. Note that some datasets are only suitable for one task or another. For example, WN18RR and HetioNet have very few relations, making it nearly impossible to sample two partitions with significantly different relations for the (E, R) task. On the other hand, FB15k-237 contains too many relations to sample multiple graphs for the (E) task without removing many edges from either the train or inference graph.
In Table 2 we show the PPR Hits@10 and $\Delta\text{SPD}$ for the inductive datasets and their original transductive dataset. We note that when multiple inference graphs exist, we take the mean across each inference graph. When possible, we also include a comparison against those inductive datasets that already exist. For example, for WN18RR in the (E) task, 4 datasets exist from [7]. As compared to the old inductive datasets, the PPR performance for the new inductive datasets is much lower. On average, the performance of PPR is 78% lower for the new inductive datasets. Furthermore, the PPR performance of the new inductive datasets is very similar to the performance on the original transductive dataset. A similar trend can be found when comparing the $\Delta\text{SPD}$. This analysis shows that newer sampling procedure can indeed sample inductive datasets that are much more similar to the original transductive graph. Furthermore, it greatly helps mitigate the PPR shortcut.
@@ 5.1 Experimental Settings
Baseline Methods: We consider prominent KGC methods including NBFNet [8], RED-GNN [11], NodePiece [19], and InGram [9]. We also consider the recent foundation model ULTRA [33].
Hyperparameters: The exact hyperparameters and their ranges differ by method due to difference in implementations. For each method, we tried to follow the recommended ranges used by the authors. For InGram [9] we tuned the learning rate in $\{1e^{-3},5e^{-4}\}$ and the number of entity layers in $\{2,4\}$. For NodePiece [19] we tuned the learning rate from $\{1e^{-3},1e^{-4}\}$ and the loss margin from $\{15,25\}$. For RED-GNN [11], we tuned the learning rate from $\{5e^{-3},5e^{-4}\}$ and the dropout from $\{0.1,0.3\}$. In their experiments, NBFNet [8] uses the same hyperparameters across all datasets. We therefore do the same.
@@ 5.2 Results
Models FB15k-237 CoDEx-M Inference 1 (27%) Inference 2 (63%) Inference 1 (10%) Inference 2 (57%) MRR Hits@10 MRR Hits@10 MRR Hits@10 MRR Hits@10 PPR 3.8 9.1 4.7 12.4 3.3 10.9 7.1 15.4 NodePiece 1.4 ± 0.1 3.0 ± 0.6 2.6 ± 0.2 4.7 ± 0.5 1.5 ± 0.2 3.1 ± 0.6 1.6 ± 0.4 2.5 ± 1.0 InGram 11.2 ± 1.9 23.8 ± 3.0 9.2 ± 0.9 20.2 ± 2.0 10.3 ± 1.2 20.4 ± 3.3 7.2 ± 4.2 15.9 ± 10.0 RED-GNN 11.8 ± 2.8 21.6 ± 5.8 19.8 ± 2.2 33.3 ± 4.2 16.8 ± 1.6 29.2 ± 2.9 14.7 ± 5.3 26.5 ± 10.4 NBFNet 15.8 ± 0.9 27.5 ± 1.8 15.9 ± 0.2 26.2 ± 0.3 27.9 ± 8.8 47.7 ± 11.8 8.5 ± 4.8 17.6 ± 10.0
Main Results: The main results for supervised methods can be found in Tables 3 and 4. We observe that on nearly every dataset, NBFNet and RED-GNN are the two best models. This indicates that conditional MPNNs [15], the class of model in which both belong to, are necessary for strong performance in inductive KGC. Interestingly, we observe that InGram struggles in the (E, R) setting. This is even true when the % of new relations is high. This run counters to the results on older inductive datasets [9] where InGram excelled over NBFNet and RED-GNN. Lastly, outside of NodePiece, each method is usually able to significantly outperform PPR.
Performance Comparison on New and Old Inductive Datasets: We compare the change in performance from the previous inductive datasets to our new datasets. Note that we only compare when there is a corresponding dataset sampled from the same transductive dataset and for the same task. For the (E) task, this includes CoDEx-M (ILPC-S [10]) and WN18RR (v1-v4 [7]). For the (E, R) task it is only FB15k-237 (25-100 [9]). For each, we compare the performance of the SOTA on the older dataset to their corresponding performance on the new datasets. This is NBFNet [8] for WN18RR and FB15k-237 and NodePiece [19] for CoDEx-M. We only compare the performance of the old SOTA as not all methods were tested on the older datasets or guranteed to be well-tuned. The results are shown in Figure 6(a). We find that performance on each of the new datasets drops significantly. This suggests that removing the shortcut has a large negative effect on the performance.
@@
Performance of ULTRA [33]: We further compare against ULTRA, a recent foundation model designed for fully inductive KGC. We evaluate ULTRA under the 0-shot setting. See Appendix E for more details on the version of ULTRA used. The results are in Figure 6(b) where for each dataset, we average the results across the different inference graphs (full results in Appendix A.2). On the (E) task, ULTRA is comparable to the supervised SOTA. However, on the (E, R) task, ULTRA significantly outperforms other methods. This suggests that ULTRA contains a much greater ability to generalize to unseen relations than existing supervised methods.
(a) Performance of SOTA on Old vs. New Datasets
(b) Results across Inf. graphs for ULTRA vs. SOTA
Figure 6: (a) Performance of SOTA on old vs. new datasets, (b) ULTRA [33] vs. supervised SOTA.
@@ 6 Conclusion
In this paper we study the problem of constructing datasets for inductive knowledge graph completion. Upon examination, we find that we can achieve competitive performance on most inductive datasets through the use of Personalized PageRank [12], which ignores the relational structure of the graph. Through our study, we uncover that this shortcut is due to how inductive datasets are created. To remedy this problem, we propose a new dataset construction process based on graph partitioning that helps mitigate the impact of the studied shortcut. We then construct new benchmark datasets using this new procedure and benchmark various methods.
@@ Appendix A Additional Results
In this section we show the performance in terms of Hits@10 for the SOTA supervised method vs. the Personalized PageRank (PPR) score. We further include the % difference in performance and which method is considered SOTA. These are shown in Tables 56, and 7 for the (E), (E, R), and transductive datasets, respectively. For an overview of SOTA performance on additional KG datasets, please see [33].
@@ B.1 Existing Datasets
We detail the statistics of all existing transductive and inductive datasets in Tables 10, 11, and 12, respectively. We further include the licenses for each in Table 13. Note that we omit YAGO3-10 [37] as [38] show that the dataset is dominated by two duplicate relations, making most triples trivial to classify. Also, we omit NELL-995 and any inductive datasets derived from it due to findings by [31] that show that most triples in the dataset are either too generic or meaningless.
@@ B.2 New Datasets
The raw data for the new datasets introduced in Section 4, can be found at https://github.com/HarryShomer/Better-Inductive-KGC/tree/master/new_data. Please see the README for more information on how to load the data.
We further detail the statistics of all the new datasets in Tables 14 and 15.
@@ Appendix D Additional Experimental Settings
We note that a lower rank, indicates better performance. In practice, some of the negative entities will be true triples observed during training. As such, following [4] we use the filtered setting and remove those entities from the ranking. Once we obtain the rank of the true entity (i.e., “how many negative entities does it have a higher probability than”) we compute two metrics – the mean reciprocal rank (MRR) and Hits@K. MRR is given by $1/\text{Rank}$ such that when Eq. 2 holds true for $\forall v\in\mathcal{V}$ it will be equal to 1. Hits@K computes whether $\text{Rank}\leq K$.
@@ Appendix F Implementation
All code used for experiments in this study can be found in the following repository – https://github.com/HarryShomer/Better-Inductive-KGC. Please see the README in the repository for more information on how to run the different experiments.

### v2

@@ 4.2 Analysis of New Datasets
In this section we analyze the new inductive datasets created following the partition-based procedure outlined in Section 4.1. We create (E) inductive datasets, from WN18RR [27], CoDEx-M [31], and HetioNet [32]. For the (E, R) datasets, we use CoDEx-M [31] and FB15k-237 [33]. Note that some datasets are only suitable for one task or another. For example, WN18RR and HetioNet have very few relations, making it nearly impossible to sample two partitions with significantly different relations for the (E, R) task. On the other hand, FB15k-237 contains too many relations to sample multiple graphs for the (E) tasks without removing many edges from either the train or inference graph.
In Table 2 we show the PPR Hits@10 and $\Delta\text{SPD}$ for the inductive datasets and their original transductive dataset. When multiple inference graphs exist, we take the mean across each inference graph. When possible, we also include a comparison against those inductive datasets that already exist. For example, for WN18RR in the (E) task, 4 datasets exist from [7]. As compared to the old inductive datasets, the PPR performance for the new inductive datasets is much lower. We find that the average PPR performance is 78% lower on the new inductive datasets as compared to the older datasets. Furthermore, the PPR performance of the new inductive datasets is very similar to the performance on the original transductive dataset. A similar trend can be found when comparing the $\Delta\text{SPD}$. This analysis shows that newer sampling procedure can indeed sample inductive datasets that are much more similar to the original transductive graph, thereby greatly mitigating the PPR shortcut.
@@ 5.1 Experimental Settings
Baseline Methods: We consider prominent KGC methods including NBFNet [8], RED-GNN [11], NodePiece [19], InGram [9], and Neural LP [13]. We also consider the recent foundation model ULTRA [34]. We omit methods that have been shown to either be prohibitively expensive (e.g.,  [7, 20, 21]) or require the use of textual information [24, 25].
@@ 5.2 Results
Main Results: The main results for the supervised methods can be found in Tables 3 and 4. We observe that on nearly every dataset, NBFNet and RED-GNN are the two best models. This indicates that conditional MPNNs [15], the class of model in which both belong to, are necessary for strong performance in inductive KGC. Interestingly, we observe that InGram struggles in the (E, R) setting. This is even true when the % of new relations is high. This runs counter to the results on older inductive datasets [9] where InGram excelled over NBFNet and RED-GNN. Lastly, we observe that Neural LP sometimes fails to converge, resulting in a near zero performance and thus causing the model to have a high performance variance.
@@
Models FB15k-237 CoDEx-M
Inference 1 (27%) Inference 2 (63%) Inference 1 (10%) Inference 2 (57%)
PPR 9.1 12.4 10.9 15.4
Neural LP 17.5 ± 9.9 22.4 ± 12.8 16.7 ± 22.8 9.8 ± 13.7
NodePiece 3.0 ± 0.6 4.7 ± 0.5 3.1 ± 0.6 2.5 ± 1.0
InGram 23.8 ± 3.0 20.2 ± 2.0 20.4 ± 3.3 15.9 ± 10.0
RED-GNN 21.6 ± 5.8 33.3 ± 4.2 29.2 ± 2.9 26.5 ± 10.4
NBFNet 27.5 ± 1.8 26.2 ± 0.3 47.7 ± 11.8 17.6 ± 10.0
Performance Comparison of PPR vs. SOTA: In our original analysis, we showed that for the older inductive datasets, the performance gap between the SOTA method of each dataset and PPR is quite small. Specifically, the mean % difference between the two was relatively small at 25-29% (see Figure 1). We now perform the same analysis on the new inductive datasets. The results are shown in Figure 6(a). We observe that PPR generally performs much worse than the SOTA method with a mean % difference of 99%. Furthermore, this is more in line with the transductive datasets which have a mean % difference of 121%. This suggests that in addition to the decrease in raw PPR performance, the relative performance of PPR also decreases on the newer datasets.
Table 5: Mean performance on Old vs. New inductive datasets by Method. Red indicates a decrease in the mean performance on the newer splits.
Models WN18RR (E) FB15k-237 (E, R)
Old Ind. New Ind. % Difference Old Ind. New Ind. % Difference
PPR 66.0 46.2 -30% 21.4 10.8 -49.5%
Neural LP 67.6 26.4 -61.0% 16.2 20.0 +23.1%
NodePiece 29.8 17.2 -42.3% 5.0 3.9 -23.0%
InGram 49.9 23.0 -53.9% 29.6 22.0 -25.7%
RED-GNN 70.6 50.3 -28.8% 25.0 27.5 +9.8%
NBFNet 72.2 52.5 -27.4% 24.8 26.9 +8.3%
Mean - - -40.6% - - -9.5%
Performance Comparison on New and Old Inductive Datasets: Previously, in Table 2, we compared both the PPR performance and $\Delta\text{SPD}$ on the new and older inductive datasets. We found that both metrics tend to be much higher on the older inductive datasets, indicating that our newer splits are effective in mitigating the shortcut. Given those results, a natural question is whether we see a similar drop in performance for neural methods? We limit our analysis to WN18RR (E) and FB15k-237 (E, R). This is either due to a lack of older datasets (i.e., CoDEX-M (E, R) and HetioNet (E)) or minimal results on the older datasets, i.e., CoDEX-M (E). For each dataset, we compute the mean performance across inference graphs. The results are shown in Table 5. We find that performance of most methods drops significantly across both datasets, with an average drop of 40.6% and 9.5% on WN18RR (E) and FB15k-237 (E, R), respectively. This suggests that removing the shortcut has a large negative effect on the performance, thereby suggesting that our new datasets are indeed harder.
@@
Performance of ULTRA [34]: We further compare against ULTRA, a recent foundation model designed for fully inductive KGC. We evaluate ULTRA under the 0-shot setting. See Appendix E for more details on the versions of ULTRA used. The results are in Figure 6(b) where for each dataset, we average the results across the different inference graphs (full results in Appendix A.2). On the (E) task, ULTRA is comparable to the supervised SOTA. However, on the (E, R) task, ULTRA significantly outperforms other methods. This suggests that ULTRA contains a much greater ability than other existing methods to generalize to unseen relations.
(a) PPR vs. SOTA Hits@10 on newer datasets
(b) Mean Hits@10 for ULTRA vs. SOTA
Figure 6: (a) Performance of PPR vs. SOTA on new inductive datasets. Note that we abbreviate CoDEx-M as CM, FB15k-237 as FB, HetioNet as HN and WN18RR as WN. Furthermore, the number represents the inference graph. (b) Mean performance of the foundation model ULTRA [34] vs. the supervised SOTA on the newer datasets.
@@ 6 Conclusion
In this paper we study the problem of constructing datasets for inductive knowledge graph completion. Upon examination, we find that we can achieve competitive performance on most inductive datasets through the use of Personalized PageRank [12], which ignores the relational structure of the graph. Through our study, we uncover that this shortcut is due to how inductive datasets are created. To remedy this problem, we propose a new dataset construction process based on graph partitioning that empirically mitigates the impact of the studied shortcut. We then construct new benchmark datasets using this new procedure and benchmark various methods. Examining the results, we observe that the relative performance decreases on most datasets and methods, indicating that the newer benchmark datasets are harder than the previous ones. For future work, we plan to explore creating inductive KG datasets that aren’t sampled from existing transductive datasets.
@@ Appendix A Additional Results
In this section we show the performance in terms of Hits@10 for the SOTA supervised method vs. the Personalized PageRank (PPR) score. We further include the % difference in performance and which method is considered SOTA. These are shown in Tables 67, and 8 for the (E), (E, R), and transductive datasets, respectively. For an overview of SOTA performance on additional KG datasets, please see [34].
@@ A.2 Performance of ULTRA
A.3 Performance of Other Potential Shortcuts
In our study, we show that PPR can achieve strong performance on most inductive KG datasets, indicating a shortcut. A natural question is whether this is true for just PPR, or if other shortcuts exist. Another bias discussed in KG literature is degree bias [38]. In their study, they show that that KG methods tend to perform better on entities with a higher degree. Specifically, they observe that what matters is the degree of the entity being predicted (they refer to this as the “tail degree”). Based on this, we study whether the tail degree can is a good predictor of performance.
We show that results on a number of inductive and transductive datasets in Table 11. We report Hits@10 for all results. We find that the tail degree often achieves a much lower performance than PPR. Specifically, for the datasets in Table  11, while the PPR performance is on average only 31% lower than the SOTA, the tail degree performance is 84% lower. This indicates the severity of the PPR shortcut is much more severe as compared to other known biases like the tail degree.
Table 11: Performance of Tail Degree [38], PPR, and SOTA method. Performance is in terms of Hits@10. We further bold the larger of the PPR and tail degree performance.
Task Dataset Tail Degree PPR SOTA
Inductive WN v1 10.4 77.1 82.6
WN v2 5.1 74.4 83.6
WN v3 10.1 45.2 58.2
WN v4 2.2 67.3 74.5
FB v1 20.5 41.2 60.7
FB v2 22.7 47.6 70.4
FB v3 16.8 43.5 67.7
FB v4 16.1 38.4 66.8
Transductive WN18RR 2.0 46.2 59.9
FB15k-237 6.0 2.7 66.6
DBPedia100k 2.9 30.2 41.8
@@ B.1 Existing Datasets
We detail the statistics of all existing transductive and inductive datasets in Tables 12, 13, and 14, respectively. We further include the licenses for each in Table 15. Note that we omit YAGO3-10 [39] as [40] show that the dataset is dominated by two duplicate relations, making most triples trivial to classify. Also, we omit NELL-995 and any inductive datasets derived from it due to findings by [31] that show that most triples in the dataset are either too generic or meaningless.
@@ B.2 New Datasets
We further detail the statistics of all the new datasets in Tables 16 and 17.
@@ Appendix D Additional Experimental Settings
Hyperparameters: Due to the different types and ranges of hyperparameters used for each method, the exact hyperparameters and their ranges differ by method. For each method, we tried to follow the recommended ranges used by the authors. For InGram [9] we tuned the learning rate in $\{1e^{-3},5e^{-4}\}$ and the number of entity layers in $\{2,4\}$. For NodePiece [19] we tuned the learning rate from $\{1e^{-3},1e^{-4}\}$ and the loss margin from $\{15,25\}$. For RED-GNN [11], we tuned the learning rate from $\{5e^{-3},5e^{-4}\}$ and the dropout from $\{0.1,0.3\}$. In their experiments, NBFNet [8] uses the same hyperparameters across all datasets. We therefore do the same. For Neural LP [13] we also use the default set of hyperparameters that is shared across datasets.
@@
We note that a lower rank, indicates better performance. In practice, some of the negative entities will be true triples observed during training. As such, following [4] we use the filtered setting and remove those entities from the ranking. Once we obtain the rank of the true entity (i.e., “how many negative entities does it have a higher probability than”) we compute the Hits@K%. Hits@Kcomputes the percentage of sample where the following inequality holds true, $\text{Rank}\leq K$.
@@ Appendix F Implementation
All code used for experiments in this study can be found in the following anonymized repository – https://github.com/HarryShomer/Better-Inductive-KGC. Please see the README in the repository for more information on how to run the different experiments.